Y= x . X^1/2 - 2/x^2 . Find the first derivative
1
Expert's answer
2019-10-15T10:23:49-0400

Solution:


We are going to find the first derivative of a function "Y= x* x^{\frac 1 2} + \frac 2 {x^{2}}"

We have a formula, Derivative of "x^{n} =" "n * x^{n-1}"


First we can rewrite the given function as Y = "x^{ 1 + \frac 1 2} + \frac 2 {x^2}" = "x^{\frac 3 2} + 2 * x^{-2}"


since, in the multiplication, If the bases are equal , we can add the power)

( If we have a term with positive exponent at the denominator, we write that at numerator with negative exponent)


Now, the derivative of the given function


Derivative of "Y = \frac 3 2 * x^{ \frac 3 2 - 1} + 2 (-2) * x^{ - 2 -1}"


= "\frac 3 2 * x^{\frac 1 2} - 4 * x^{ -3}"


Answer: Derivative of Y ="\frac 3 2 * x^{\frac 1 2} - \frac 4 {x^{3}}"